# The vertical structure of debris discs and the role of disc gravity

### v1

@@ 4 Secular behavior: analytical results
where $\Delta\Omega=\Omega_{d}-\Omega_{p}$ stays in the range $[-\pi,\pi]$. In Equation (18), the forced planetesimal inclination $I_{d,\rm{forced}}$ is:
@@ 4.1 Dynamical regimes
In Equation (4.1), we have assumed $a_{p}/a_{d}\ll 1$, so that $b_{3/2}^{(1)}(\alpha)\approx 3\alpha$, and $C\equiv(1-\delta^{p-2})/(\delta^{p+1}-1)$. Note that for $\delta=a_{\rm out}/a_{\rm in}\gg 1$, one has $C\delta^{3}\rightarrow\delta^{2-p}$ for $20<p<$ 2, i.e., for disks that contain more mass in the outer parts than in the inner parts (Equation 1). Accordingly, the ratio $A_{d,p}/A_{p}$ is positive and, more importantly, an increasing function of $a_{d}/a_{\rm in}$ for all astrophysically motivated values of $p$; see also Figure 2. Note that $A_{d,p}/A_{p}$ depends on the planet and disk masses only through their ratio, $M_{d}/m_{p}$, and is independent of $M_{c}$.
@@ 4.1.1 Secular inclination resonance
A 00 $1M_{J}$ $20$ au Planet-dominated: entire disk $10^{\circ}$infty $139.6$ Myr $T_{{\rm sec}}(80~{}{\rm au})\approx 14.2$ Myr
@@ 4.2.1 Secular evolution of the planetary orbit
Note that for a massless disk, $A_{d,p}=0$ and the planetary orbit does not evolve, that is, $I_{p}(t)=I_{p}(0)$ and $\Omega_{p}(t)=\Omega_{p}(0)$. Before moving on, we emphasize that when the full back-reaction of the disk is taken into account, the planetary inclination may evolve over time; see Appendix B and Section 6.4 for further discussion.
@@ 4.2.2 Secular evolution of the planetesimal orbits
In conjunction with planetary precession, planetesimal orbits undergo both nodal precession and inclination oscillations. The inclinations oscillate between the initial value of 00 and a maximum of $I_{m}(a_{d})=2|I_{d,\rm{forced}}|=2\langle I_{d}(t)\rangle_{t}$; see Equations (18)–(20) and Figure 5. The oscillation period is
$T_{\rm sec}(a_{d})=\frac{2\pi}{|A_{p}(a_{d})-A_{d,p}|},$ (28)
@@
irrespective of their semimajor axes; see Equation (20) and Figure 5(a). Thus, when time-averaged, each planetesimal lies within the plane of the planetary orbit, $\langle I_{d}(t)\rangle_{t}^{{\rm n/disk}}=I_{p}(0)$, and the disk’s angular momentum vector is aligned with that of the planet. This corresponds to the well-studied limiting case of a massless disk (e.g., Dawson et al., 2011). In the course of oscillations, planetesimal inclinations are maximized when their orbits are nodally aligned with the planet, $\Delta\Omega(t)=0$, and, since the forced inclination is positive, $\Delta\Omega(t)$ remains within the range $[-\pi/2,\pi/2]$ – see also Figure 5(a). Finally, the oscillations occur with a period of:
@@
see Equations (20) and (4.1). Comparing $I_{m}^{{\rm disk}}$ with $I_{m}^{{\rm n/disk}}$, two key differences emerge: first, planetesimal inclinations in the disk-dominated regime decrease with distance ($\propto a_{d}^{-7/2}$); second, their amplitudes can be much smaller than those in the planet-dominated regime, with $I_{m}^{{\rm disk}}\lesssim I_{m}^{{\rm n/disk}}$ – see also Figures 3 and 5(b), (c). Relatedly, when time-averaged, the angular momentum vectors of planetesimals will be misaligned with that of the planet, with the degree of misalignment increasing as the planet-planetesimal separation grows. Finally, since the forced inclination is negative in this regime, $\Delta\Omega(t)$ oscillations are confined to the range ±$\pm[\pi/2,\pi]$, and inclinations are maximized when planetesimals are nodally anti-aligned with the planet, $\Delta\Omega=\pi$; see also Figures 5(b), (c). Such oscillations occur with a period of (Equation (28)):
@@
which, unlike $T_{{\rm sec}}^{{\rm n/disk}}$ (Equation 30), is independent of planetesimal semimajor axis (see also Equation (27)).
@@
while their orbits remain locked to the planetary node such that $\Delta\Omega(t)=\pi/2$; see also Figure 5(b). Thus, the timescale $\tau_{\rm res}$ for the inclinations to grow from their initial value of 00 to $\pi/2$, i.e., $I_{d}(t)/90^{\circ}=t/\tau_{\rm res}$, can be approximated as:
@@ 5 Evolution of the disk Morphology
Figure 6: Snapshots showing the time evolution of the surface density $\Sigma_{d}$ of the debris disk in Model A (i.e., $M_{d}=0$; Table 1), viewed from three orientations: face-on (top row), end-on (center row), and edge-on (bottom row). The snapshots correspond to the same times $t$ as those in Figure 5(a), as indicated above each column (left to right). In each panel, the stellar position is marked by the yellow star, while the planet’s orbit and its (ascending) nodal position are shown by the white line and red circle, respectively. To highlight fine details, surface densities are plotted on a logarithmic scale (as indicated by the color bar), with all panels normalized by $\Sigma_{d}^{t=0}(a_{\rm in})$. The density maps derive from $N=5,000$ parent planetesimals, each with $N_{np}=10,000$ particles per orbit, populating the $3$-D plane divided into equally sized cubic cells, each with a volume of $(0.3~{}{\mathrm{au}})^{3}$. Calculations are done using the analytically derived distribution of orbital elements (rather than direct $N$-body simulations; see Appendix D.1 for technical details). It is evident that at early times (panels (i)), the planet launches a leading bending wave at the disk’s inner edge, $a_{\rm in}$, which appears as a warp when viewed edge-on and as a double-armed spiral when viewed face-on. The location of the warp, where the disk reaches its maximum vertical height from the $(X,Y)$ plane, moves outward over time (panels (i)–(iv)). The warp feature fades away upon reaching the disk’s outer edge $a_{\rm out}$ (panel (iv)); which occurs after half a secular precession period at $a_{\rm out}$, i.e., $t=0.5T_{\rm sec}(a_{\rm out})$. Beyond this time (panels (v), (vi)), the entire disk becomes fully inclined, forming a thick, box-like structure characterized by an “X”-shaped density pattern that is symmetrical about the planetary orbital plane. In other words, the disk and the planet are aligned. At late times, the disk’s aspect ratio is distance-independent, given by $\mathcal{H}\approx I_{p}(0)$. See the text (Section 5.1) for further details.
@@
To facilitate the presentation, we consider the same three model systems analyzed in Section 4, the parameters and relevant secular timescales of which are summarized in Table 1. We begin by considering Model A (Section 5.1), in which the disk is massless ($M_{d}=0$). This serves as a reference case for comparison with the results for massive disks, starting with $M_{d}/m_{p}\lesssim 1$ (Model B; Section 5.2) and then $M_{d}/m_{p}\gtrsim 1$ (Model C; Section 5.3). To this end, we convert the distributions of orbital elements shown in Figure 5 – which, we recall, were obtained using the analytical solutions given by Equations (16)–(20) – into distributions of disk surface density $\Sigma_{d}(t)$. To provide a comprehensive view, density maps are then projected onto three planes: namely, (i) $X-Y$ (face-on view, with the observer looking along the $Z$-axis), (ii) $X-Z$ (end-on view, with the observer looking along the $Y$-axis), and (iii) $Y-Z$ (edge-on view, with the observer looking along the $X$-axis). Technical details of this procedure are provided in Appendix D.1. For clarity, we note that the disk initially lies in the $X-Y$ plane ($Z=0$), with the positive $X$-axis pointing towards the planet’s initial ascending node. We close this section by examining the effects of varying system parameters relative to their fiducial values (Section 5.4).
@@ 5.1 Massless Disks ($M_{d}/m_{p}=0$)
Stage 1 $[0\leq t<0.5T_{\rm sec}^{{\rm n/disk}}(a_{\rm out})]$: At early times, the disk rapidly departs from its initial axisymmetric, razor-thin state by developing a spiral bending wave (Fig. 6(i)). The bending wave, launched at the disk’s inner edge ($a_{\rm in}=30$ au), travels outward as a leading pattern, i.e., rotating in the same direction as the planetesimals’ motion. In the top-down images, the bending wave appears as a double-armed spiral, while in the edge-on images, it manifests as vertical oscillations (more on this later). The leading nature of the wave follows from the fact that $d\Omega_{d}/da_{d}>0$ at all times (see, e.g., Figure 5(a)). The wave’s propagation rate is determined by the planet’s mass and semimajor axis, causing it to propagate more slowly as it extends to larger radii. For instance, the wave travels from $a_{\rm in}$ to $\sim 80$ au in $\approx 7$ Myr, and it takes an additional $\approx 24$ Myr to extend to $\sim 120$ au (Figs. 6(i)–(iii)). This is because $|A_{p}(a_{d})|\propto a_{d}^{-7/2}$, rendering $T_{\rm sec}^{{\rm n/disk}}(a_{d})$ longer at larger distances; see Eq. (30).
As the waves travel outward, they wrap around the star, becoming more tightly wound closer to the planet than at larger distances; see, e.g., Figs. 6(i)–(iv). The outermost portion of the bending wave is associated with planetesimals at $a_{d}=a_{\rm warp}$ that are in the midst of their first oscillation cycle about the planetary orbit: that is, attaining their maximum inclinations for the first time with $I_{m}(a_{d})=2I_{p}(0)$ and $\Delta\Omega=0$; see Figure 5(a). Interior to this region, the waves are tightly wound, making them difficult to discern. This is because planetesimals have completed more than one precession cycle, and their inclinations and longitudes of ascending nodes are relatively well-mixed within the ranges $[0,2I_{p}(0)]$ and $[-\pi/2,\pi/2]$, respectively (Figure 5(a)). Nevertheless, some windings remain identifiable, particularly just inside the outermost portion of the waves. In the top-down images, these appear as a double-armed spiral: a structure that is closely reproduced by the nodal positions of planetesimal orbits (both ascending and descending, i.e., $[X,Y,Z]=\pm a_{d}[\cos\Omega_{d},\sin\Omega_{d},0]$; Equation (71)), along with overdense regions due to planetesimals that have not yet phase-mixed. A complimentary view of this can be seen in the vertical oscillations apparent in e.g. the edge-on images.
At $a_{d}=a_{\rm warp}$, the disk attains its maximum vertical extent $|z_{\rm warp}|$ from the initial midplane, featuring a warped structure, as best observed in the edge-on images (Figs. 6(i)–(iv)). The maximum height $|z_{\rm warp}|$ from the $X-Y$ plane is:
@@
Stage 2 $[t\geq 0.5T_{\rm sec}^{{\rm n/disk}}(a_{\rm out})]$: The warp, however, is a transient phenomenon. Indeed, at later times, i.e., when $t=0.5T_{\rm sec}^{{\rm n/disk}}(a_{\rm out})$, the warp location reaches the disk’s outer edge, $a_{\rm warp}=a_{\rm out}$ (Figure 6(iv)). Subsequently, when all planetesimals have completed at least one full precession cycle, i.e., when $t\gtrsim T_{\rm sec}^{{\rm n/disk}}(a_{\rm out})$, the warp structure disappears entirely (Figures 6(v), (vi)).
@@
When this occurs, the disk settles into a quasi-equilibrium state, characterized by a fully inclined, thick, box-like structure that is symmetric about the planetary orbit. This is because, at such late times, planetesimal orbits become relatively well-mixed such that $0\leq I_{d}(t)\leq 2I_{p}(0)$ and $|\Delta\Omega|\leq\pi/2$ at all $a_{d}$ (Figure 5(a)). This implies that the total angular momentum vector of the disk is aligned with that of the planet, both tilted by $I_{p}(0)$ relative to the initial disk midplane. This is illustrated by the blue curve in Figure 7, which shows the time evolution of $\theta_{{\rm al}}$, the angle between the planet’s and the disk’s angular momentum vectors (Appendix D.3). It is evident that $\theta_{{\rm al}}$ decreases from an initial value of $I_{p}(0)$ to 00 over time. Relatedly, the disk acquires a vertical scale height, $\mathfrak{h}_{d}$, or equivalently an aspect ratio, $\mathcal{H}\equiv\mathfrak{h}_{d}/R$, set by the planet. Namely, the scale height increases linearly with radius, $\mathfrak{h}_{d}\propto R$, so that the aspect ratio is a distance-independent constant given by $\mathcal{H}\approx I_{p}(0)$. Figure 8 illustrates the radial dependence of the aspect ratio as a function of time, as measured from the density maps (see Appendix D.2 for technical details).
@@
Figure 9: Similar to Figure 6, but for Model B (Table 1). Model B is identical to Model A, except that its debris disk has a non-zero mass of $M_{d}/m_{p}=0.1$. Density maps are generated using the same method (Appendix D.1) and parameters as in Figure 6, including grid cell volume, $N$, $N_{np}$, and the normalization constant. The snapshots (left to right) correspond to the same times $t$ as in Figure 5(b). All other notations and symbols are the same as in Figures 5(b) and 6. It is evident that the disk’s back-reaction on the planet significantly affects its resultant morphology. At early times (panels (i),(ii)), a bending wave launched at the disk’s inner edge propagates outward; this appears as a warp (two-armed spiral) when viewed edge-on (top-down). The region interior to the warp is tilted and corotates with the planetary node, while remaining aligned and symmetric around the planet’s orbit. By the time the planet completes one precession cycle (panel (iii)), the warp extends out to $a_{\rm res}\approx 80.2$ au, where it stalls at later times. The disk’s inner part develops a curly “X”-pattern, whose out-of-plane arm does not connect smoothly with the outer parts, thus the appearance of a “broken” disk. Beyond this time (panels (iv)–(vi)), the disk structure remains practically invariant, coprecessing with the planetary node. In terms of angular momentum, the disk maintains the initial misalignment with the planet, while its vertical aspect ratio develops a radial dependence (unlike when $M_{d}=0$). See the text (Section 5.2) for further details.

### v2

@@ 4 Secular behavior: simplified analytical results
where $\Delta\Omega=\Omega_{d}-\Omega_{p}$ stays in the range $\pi,\pi]$. In Equation (18), the forced planetesimal inclination $I_{d,\rm{forced}}$ is:
@@ 4.1 Dynamical regimes
In Equation (4.1), we have assumed $a_{p}/a_{d}\ll 1$, so that $b_{3/2}^{(1)}(\alpha)\approx 3\alpha$, and $C\equiv(1-\delta^{p-2})/(\delta^{p+1}-1)$. Note that for $\delta=a_{\rm out}/a_{\rm in}\gg 1$, one has $C\delta^{3}\rightarrow\delta^{2-p}$ for $0<p<2$, i.e., for disks that contain more mass in the outer parts than in the inner parts (Equation 1). Accordingly, the ratio $A_{d,p}/A_{p}$ is positive and, more importantly, an increasing function of $a_{d}/a_{\rm in}$ for all astrophysically motivated values of $p$; see also Figure 2. Note that $A_{d,p}/A_{p}$ depends on the planet and disk masses only through their ratio, $M_{d}/m_{p}$, and is independent of $M_{c}$.
@@ 4.1.1 Secular inclination resonance
A 0 $1M_{J}$ 2020 au Planet-dominated: entire disk $10^{\circ}$ $\infty$ $139.6$ Myr $T_{{\rm sec}}(80~{\rm au})\approx 14.2$ Myr
@@ 4.2.1 Secular evolution of the planetary orbit
Note that for a massless disk, $A_{d,p}=0$ and the planetary orbit does not evolve, that is, $I_{p}(t)=I_{p}(0)$ and $\Omega_{p}(t)=\Omega_{p}(0)$. Before moving on, we emphasize that when the full back-reaction of the disk is taken into account, the planetary inclination may evolve over time (e.g., decay exponentially) and the planet’s nodal precession period may deviate by a factor of a few from that given by Equation (27); see Appendix B and Section 6.4 for further discussion.
@@ 4.2.2 Secular evolution of the planetesimal orbits
In conjunction with planetary precession, planetesimal orbits undergo both nodal precession and inclination oscillations. The inclinations oscillate between the initial value of 0 and a maximum of $I_{m}(a_{d})=2|I_{d,\rm{forced}}|=2\langle I_{d}(t)\rangle_{t}$; see Equations (18)–(20) and Figure 5. The oscillation period is
$T_{\rm sec}(a_{d})=\frac{2\pi}{|A(a_{d})|}\equiv\frac{2\pi}{|A_{p}(a_{d})-A_{d,p}|},$ (28)
@@
irrespective of their semimajor axes; see Equation (20) and Figure 5(a). Thus, when time-averaged, each planetesimal lies within the plane of the planetary orbit, $\langle I_{d}(t)\rangle_{t}^{{\rm n/disk}}=I_{p}(0)$, and the disk’s angular momentum vector is aligned with that of the planet. This corresponds to the well-studied limiting case of a massless disk (e.g., Dawson et al., 2011). In the course of oscillations, planetesimal inclinations are maximized when their orbits are nodally aligned with the planet, $\Delta\Omega(t)=0$, and, since the forced inclination is positive, $\Delta\Omega(t)$ remains within the range $\pi/2,\pi/2]$ – see also Figure 5(a). Finally, the oscillations occur with a period of:
@@
see Equations (20) and (4.1). Comparing $I_{m}^{{\rm disk}}$ with $I_{m}^{{\rm n/disk}}$, two key differences emerge: first, planetesimal inclinations in the disk-dominated regime decrease with distance ($\propto a_{d}^{-7/2}$); second, their amplitudes can be much smaller than those in the planet-dominated regime, with $I_{m}^{{\rm disk}}\lesssim I_{m}^{{\rm n/disk}}$ – see also Figures 3 and 5(b), (c). Relatedly, when time-averaged, the angular momentum vectors of planetesimals will be misaligned with that of the planet, with the degree of misalignment increasing as the planet-planetesimal separation grows. Finally, since the forced inclination is negative in this regime, $\Delta\Omega(t)$ oscillations are confined to the range $\pm[\pi/2,\pi]$, and inclinations are maximized when planetesimals are nodally anti-aligned with the planet, $\Delta\Omega=\pi$; see also Figures 5(b), (c). Such oscillations occur with a period of (Equation (28)):
@@
where we have assumed that $|A_{d,p}/A_{p}|\gg 1$. Note that, unlike $T_{{\rm sec}}^{{\rm n/disk}}$ (Equation 30), $T_{{\rm sec}}^{{\rm disk}}$ is independent of planetesimal semimajor axis (see also Equation (27)). We clarify, however, that this does not necessarily imply that the planetesimal orbits remain strictly nodally aligned, as $A_{p}$ is small (compared to $A_{d,p}$) but nonzero in the disc-dominated regime. This subtle contribution accounts for the differential evolution of $\Delta\Omega(a_{d})$ observed in the relevant regions of Figures 5(b) and (c).
@@
while their orbits remain locked to the planetary node such that $\Delta\Omega(t)=\pi/2$; see also Figure 5(b). Thus, the timescale $\tau_{\rm res}$ for the inclinations to grow from their initial value of 0 to $\pi/2$, i.e., $I_{d}(t)/90^{\circ}=t/\tau_{\rm res}$, can be approximated as:
@@ 5 Evolution of the disk Morphology
Figure 6: Snapshots showing the time evolution of the surface density $\Sigma_{d}$ of the debris disk in Model A (i.e., $M_{d}=0$; Table 1), viewed from three orientations: face-on (top row), end-on (center row), and edge-on (bottom row). The snapshots correspond to the same times $t$ as those in Figure 5(a), as indicated above each column (left to right). In each panel, the stellar position is marked by the yellow star, while the planet’s orbit and its (ascending) nodal position are shown by the white line and red circle, respectively. To highlight fine details, surface densities are plotted on a logarithmic scale (as indicated by the color bar), with all panels normalized by $\Sigma_{d}^{t=0}(a_{\rm in})$. The density maps derive from $N=5,000$ parent planetesimals, each with $N_{np}=10,000$ particles per orbit, populating the 33-D volume divided into equally sized cubic cells, each with a volume of $(0.3~{\mathrm{au}})^{3}$. Calculations are done using the analytically derived distribution of orbital elements (rather than direct $N$-body simulations; see Appendix D.1 for technical details). It is evident that at early times (panels (i)), the planet launches a leading bending wave at the disk’s inner edge, $a_{\rm in}$, which appears as a warp when viewed edge-on and as a double-armed spiral when viewed face-on. The location of the warp, where the disk reaches its maximum vertical height from the $(X,Y)$ plane, moves outward over time (panels (i)–(iv)). The warp feature fades away upon reaching the disk’s outer edge $a_{\rm out}$ (panel (iv)); which occurs after half a secular precession period at $a_{\rm out}$, i.e., $t=0.5T_{\rm sec}(a_{\rm out})$. Beyond this time (panels (v), (vi)), the entire disk becomes fully inclined, forming a thick, box-like structure characterized by an “X”-shaped density pattern that is symmetrical about the planetary orbital plane. In other words, the disk and the planet are aligned. At late times, the disk’s aspect ratio is distance-independent, given by $\mathcal{H}\approx I_{p}(0)$. See the text (Section 5.1) for further details.
@@
To facilitate the presentation, we consider the same three model systems analyzed in Section 4, the parameters and relevant secular timescales of which are summarized in Table 1. We begin by considering Model A (Section 5.1), in which the disk is massless ($M_{d}=0$). This serves as a reference case for comparison with the results for massive disks, starting with $M_{d}/m_{p}\lesssim 1$ (Model B; Section 5.2) and then $M_{d}/m_{p}\gtrsim 1$ (Model C; Section 5.3). To this end, we convert the distributions of orbital elements shown in Figure 5 – which, we recall, were obtained using the simplified analytical solutions given by Equations (16)–(20) – into distributions of disk surface density $\Sigma_{d}(t)$. To provide a comprehensive view, density maps are then projected onto three planes: namely, (i) XYX-Y (face-on view, with the observer looking along the $Z$-axis), (ii) XZX-Z (end-on view, with the observer looking along the $Y$-axis), and (iii) YZY-Z (edge-on view, with the observer looking along the $X$-axis). Technical details of this procedure are provided in Appendix D.1. For clarity, we note that the disk initially lies in the XYX-Y plane ($Z=0$), with the positive $X$-axis pointing towards the planet’s initial ascending node. We close this section by examining the effects of varying system parameters relative to their fiducial values (Section 5.4).
From here on, unless otherwise specified, we define the disk’s scale height, $\mathfrak{h}_{d}$, as the square root of the mass-weighted second moment of the planetesimal’s vertical distribution about their mean height (see Appendix D.2 for details). The corresponding local aspect ratio is then given by $\mathcal{H}(R)=\mathfrak{h}_{d}/R$, where $R=\sqrt{X^{2}+Y^{2}}$ is the cylindrical radius. This definition is adopted because the shape of the vertical density is not known a priori.
@@ 5.1 Massless Disks ($M_{d}/m_{p}=0$)
Stage 1 $0\leq t<0.5T_{\rm sec}^{{\rm n/disk}}(a_{\rm out})]$: At early times, the disk rapidly departs from its initial axisymmetric, razor-thin state by developing a spiral bending wave (Fig. 6(i)). The bending wave, launched at the disk’s inner edge ($a_{\rm in}=30$ au), travels outward as a leading pattern, i.e., rotating in the same direction as the planetesimals’ motion. In the top-down images, the bending wave appears as a double-armed spiral, while in the edge-on images, it manifests as vertical oscillations (more on this later). The leading nature of the wave follows from the fact that $d\Omega_{d}/da_{d}>0$ at all times (see, e.g., Figure 5(a)). The wave’s propagation rate is determined by the planet’s mass and semimajor axis, causing it to propagate more slowly as it extends to larger radii. For instance, the wave travels from $a_{\rm in}$ to $\sim 80$ au in $\approx 7$ Myr, and it takes an additional $\approx 24$ Myr to extend to $\sim 120$ au (Figs. 6(i)–(iii)). This is because $|A_{p}(a_{d})|\propto a_{d}^{-7/2}$, rendering $T_{\rm sec}^{{\rm n/disk}}(a_{d})$ longer at larger distances; see Eq. (30).
As the waves travel outward, they wrap around the star, becoming more tightly wound closer to the planet than at larger distances; see, e.g., Figs. 6(i)–(iv). The outermost portion of the bending wave is associated with planetesimals at $a_{d}=a_{\rm warp}$ that are in the midst of their first oscillation cycle about the planetary orbit: that is, attaining their maximum inclinations for the first time with $I_{m}(a_{d})=2I_{p}(0)$ and $\Delta\Omega=0$; see Figure 5(a). Interior to this region, the waves are tightly wound, making them difficult to discern. This is because planetesimals have completed more than one precession cycle, and their inclinations and longitudes of ascending nodes are relatively well-mixed within the ranges $0,2I_{p}(0)]$ and $\pi/2,\pi/2]$, respectively (Figure 5(a)). Nevertheless, some windings remain identifiable, particularly just inside the outermost portion of the waves. In the top-down images, these appear as a double-armed spiral: a structure that is closely reproduced by the nodal positions of planetesimal orbits (both ascending and descending, i.e., $X,Y,Z]=\pm a_{d}[\cos\Omega_{d},\sin\Omega_{d},0]$; Equation (71)), along with overdense regions due to planetesimals that have not yet phase-mixed. A complimentary view of this can be seen in the vertical oscillations apparent in e.g. the edge-on images.
At $a_{d}=a_{\rm warp}$, the disk attains its maximum vertical extent $|z_{\rm warp}|$ from the initial midplane, featuring a warped structure, as best observed in the edge-on images (Figs. 6(i)–(iv)). The maximum height $|z_{\rm warp}|$ from the XYX-Y plane is:
@@
Stage 2 $t\geq 0.5T_{\rm sec}^{{\rm n/disk}}(a_{\rm out})]$: The warp, however, is a transient phenomenon. Indeed, at later times, i.e., when $t=0.5T_{\rm sec}^{{\rm n/disk}}(a_{\rm out})$, the warp location reaches the disk’s outer edge, $a_{\rm warp}=a_{\rm out}$ (Figure 6(iv)). Subsequently, when all planetesimals have completed at least one full precession cycle, i.e., when $t\gtrsim T_{\rm sec}^{{\rm n/disk}}(a_{\rm out})$, the warp structure disappears entirely (Figures 6(v), (vi)).
@@
When this occurs, the disk settles into a quasi-equilibrium state, characterized by a fully inclined, thick, box-like structure that is symmetric about the planetary orbit. This is because, at such late times, planetesimal orbits become relatively well-mixed such that $0\leq I_{d}(t)\leq 2I_{p}(0)$ and $|\Delta\Omega|\leq\pi/2$ at all $a_{d}$ (Figure 5(a)). This implies that the total angular momentum vector of the disk is aligned with that of the planet, both tilted by $I_{p}(0)$ relative to the initial disk midplane. This is illustrated by the blue curve in Figure 7, which shows the time evolution of $\theta_{{\rm al}}$, the angle between the planet’s and the disk’s angular momentum vectors (Appendix D.3). It is evident that $\theta_{{\rm al}}$ decreases from an initial value of $I_{p}(0)$ to 0 over time. Relatedly, the disk acquires a vertical scale height, $\mathfrak{h}_{d}$, or equivalently an aspect ratio, $\mathcal{H}\equiv\mathfrak{h}_{d}/R$, set by the planet. Namely, the scale height increases linearly with radius, $\mathfrak{h}_{d}\propto R$, so that the aspect ratio is a distance-independent constant given by $\mathcal{H}\approx I_{p}(0)$. Figure 8 illustrates the radial dependence of the aspect ratio as a function of time, as measured from the density maps (see Appendix D.2 for technical details).
@@
Figure 9: Similar to Figure 6, but for Model B (Table 1). Model B is identical to Model A, except that its debris disk has a non-zero mass of $M_{d}/m_{p}=0.1$. Density maps are generated using the same method (Appendix D.1) and parameters as in Figure 6, including grid cell volume, NN, $N_{np}$, and the normalization constant. The snapshots (left to right) correspond to the same times $t$ as in Figure 5(b). All other notations and symbols are the same as in Figures 5(b) and 6. It is evident that the disk’s back-reaction on the planet significantly affects its resultant morphology. At early times (panels (i),(ii)), a bending wave launched at the disk’s inner edge propagates outward; this appears as a warp (two-armed spiral) when viewed edge-on (top-down). The region interior to the warp is tilted and corotates with the planetary node, while remaining aligned and symmetric around the planet’s orbit. By the time the planet completes one precession cycle (panel (iii)), the warp extends out to $a_{\rm res}\approx 80.2$ au, where it stalls at later times. The disk’s inner part develops a curly “X”-pattern, whose out-of-plane arm does not connect smoothly with the outer parts, thus the appearance of a “broken” disk. Beyond this time (panels (iv)–(vi)), the disk structure remains practically invariant, coprecessing with the planetary node. In terms of angular momentum, the disk maintains the initial misalignment with the planet, while its vertical aspect ratio develops a radial dependence (unlike when $M_{d}=0$). See the text (Section 5.2) for further details.